Given the following variables in the open economy aggregate expenditure model, autonomous consumption (co)=200, autonomous investment (Io)=200, government spending (G0)=100, eport spending (Xo)= 100, autonomous import spending (Mo)=100, taxes (Tp)=0, marginal propensity to invest (i1)=0.1, and marginal propensity to import (m1) =0.15 calculate the equilibrium level of income for he open economy aggregate expediture model.
1
Expert's answer
2019-08-07T09:17:56-0400

C0 = 200, I0 = 200, G0 = 100, X0 = 100, M0 = 100, T = 0, c1 = 0.75, i1 = 0.1, m1 = 0.15.


The equilibrium level of income is given by

Y = C + I + G + NX = C0 + c1(Y - T) + I0 + i1*Y + G0 + X0 - (M0 + m1*Y),

Y = 200 + 0.75(Y - 0) + 200 + 0.1Y + 100 + 100 - (100 + 0.15Y),

Y = 500 + 0.7Y,

0.3Y = 1650,

Y = 5500